# Exploring the Potential of Multimodal LLM with Knowledge-Intensive Multimodal ASR

### v1

@@ Abstract
Recent advancements in multimodal large language models (MLLMs) have made significant progress in integrating information across various modalities, yet real-world applications in educational and scientific domains remain challenging. This paper introduces the Multimodal Scientific ASR (MS-ASR) task, which focuses on transcribing scientific conference videos by leveraging visual information from slides to enhance the accuracy of technical terminologies. Realized that traditional metrics like WER fall short in assessing performance accurately, prompting the proposal of severity-aware WER (SWER) that considers the content type and severity of ASR errors. We propose the Scientific Vision Augmented ASR (SciVASR) framework as a baseline method, enabling MLLMs to improve transcript quality through post-editing. Evaluations of state-of-the-art MLLMs, including GPT-4o, show a 45% improvement over speech-only baselines, highlighting the importance of multimodal information integration.
@@
Minghan Wang1, Yuxia Wang2, Thuy-Trang Vu, Ehsan Shareghi1, Gholamreza Haffari1 1Department of Data Science & AI, Monash University  2MBZUAI {minghan.wang,trang.vu1,ehsan.shareghi,gholamreza.haffari}@monash.edu yuxia.wang@mbzuai.ac.ae
@@ 1 Introduction
One such challenging task is multimodal automatic speech recognition (ASR) from knowledge-intensive presentation videos, which involves not only transcribing spoken content but also understanding and integrating visual information from images or videos (Oneata and Cucu, 2022; Chang et al., 2023). Presentation videos, often found in scientific conferences or educational contexts, contain dense, knowledge-rich content spanning various topics and concepts. Effective recognition and comprehension of such content are crucial for applications such as automatic transcript generation (Angrave et al., 2020), lecture summarization (Benedetto et al., 2024) and accessibility tools for researchers and learners (McCarron, 2021). ASR on such videos presents several challenges, including keeping pace with the rapid evolution of technical domain knowledge and handling the variability in speaker accents, proficiency in English, and speech fluency (Cettolo et al., 2004). Mitigating these challenges requires ASR models to effectively extract and incorporate information from the speech and presentation slides in the video recordings (see Figure 1).
@@ 3.1 Speech Recognition
Figure 2: The architecture of AcaVASR. Baseline ASR model transcribes a presentation recording into a transcript. Video segmenter splits video frames into a sequence of scenes with each presenting a slide. Slide analyzer applies multimodal LLMs to extract key information from scene images into visual contexts represented by text. Post-editor instruct a text-LLM to leverage visual contexts to refine the ASR transcript.
@@ 4 Experiments
Table 2: This table presents the full settings of our experiment showing the result in WER and Term-Recall (TR). Videos with $\star$ are from the test set. For clarity, we show the relative improvements compared to the Whisper-base with red and green distinguishing worse or better performance. Best results are highlighted with bold.
@@ 4.1 Models
ASR Models We experiment with Whisper-base and Whisper-large-v3 (Radford et al., 2023) as speech recognition models. To challenge the effectiveness of our proposed method, we use the inferior Whisper-base as the baseline ASR whose outputs demands more complex refinement.
Vision-LLMs We evaluate both open-sourced model CogAgent-VQA (Hong et al., 2023; Wang et al., 2023c) given the strengths in GUI and textual-related VQA tasks, and closed-sourced model GPT-4o (OpenAI et al., 2024).
Text-LLMs Based on the observation that models with less than 13B parameters often fail to follow complex instructions and exhibit severe hallucinations, we apply open-sourced Llama3-70B-Instruct (AI@Meta, 2024) and closed-sourced GPT-4o in our experiments for effective error correction.
@@ 4.3 Results and Analysis
Why does post-editing increase WER? Despite the significant improvements in term recall, WER doesn’t reveal consistent improvements. Open-sourced PE methods even increase the WER in 6 out of 10 videos. The discrepancy and the counter-intuition motivate us to investigate why the improvements aren’t reflected on WER. By analyzing the model outputs, we find that while post-editing accurately corrects a certain amount (roughly 40%) of explicit homophone errors, WER still judges them as errors due to the tiny differences with ground truth text in spelling, e.g., present vs. presents, and fine-tuning vs. finetuning. Additionally, acceptable minor rephrasing caused by the post-editing also increases WER, despite the improvements in fluency. These negligible errors are penalized similarly in WER compared with severe errors like recognizing BERT as birds, making WER fail to reflect the true quality of the transcript, significantly deviating from human judgments as shown in Figure 3. Post-edited transcripts are consistently voted to be better than the baseline transcript, and Vision-PE is more preferred than Text-PE results. To address this, we propose a severity-aware ASR evaluation protocol in §5.
Figure 3: The win, tie and lose ratio between different settings by human annotations, based on Whisper-base and AcaVASR in open-source setup.
@@ 5.1 Breakdown of WER
WER is computed by counting the number of required operations (insertion, omission, substitution – $I,O,S$) to transform the hypothesis into the reference and then normalizing it by the reference length $N$. Each operation relates to a mismatch in the hypothesis. Formally, WER is calculated as $\text{WER}=\frac{I+O+S}{N}$ where it assigns equal weights to all operations, without considering the varying importance and the semantic impact of the different types of mismatched content. This makes the phenomenon that a transcript with ten errors on general words like “the” or “a” has the same WER as a transcript with ten errors in terminologies. We mitigate it by re-weighting mismatches according to the content type and error severity.
@@
Content Type Catagory
@@ 5.3 Severity-Aware WER
where $N$ is the reference length, $K$ is the number of mismatches which is same as the total number of operations required in WER, and $w_{i}$ is the weight corresponding to the severity level of each mismatch. In our experiment, we simply define the mapping of severity levels to weights as $w_{critical}$ = 1.0, $w_{minor}$ = 0.6, and $w_{ok}$ = 0.2.333The rationale is to down-weight less severe mismatches and place greater emphasis on critical errors. We also tried the setting of (1.0, 0.5, 0.1) and found that (1.0, 0.6, 0.2) gives a better correlation with human evaluations on the dev set.
ASR Baselines
AcaVASR with Open-sourced LLMs (Llama3-70B-Instruct + CogAgent-VQA)
AcaVASR with Proprietary LLMs (GPT-4o)
@@ 5.4 How reliable is LLM evaluation?
We prompt an LLM to label each mismatch with its content type category and severity level as shown in Figure 7 (see detailed prompt in Figure 6 in Appendix C). GPT-4o is used in our experiments.
@@ 6 Re-Evaluating SciVASR
Compared with ASR baselines, text-only post-editing without visual information reduces both SWER and the content-wise severity scores to some extent. SWER decreases from 5.0 to 4.76 in open-sourced setting and 4.08 in close-sourced setting, indicating that parameterized knowledge in text-LLM is helpful in correcting potential errors.
Visual context are critical for correcting termonology, named entity and numbers
Incorporating visual information brings a significant reduction in the severity scores for TERM (-152), NE (-38) and NUM (-5) using GPT-4o, contrasting to Whisper-base. SWER reaches 2.75 from 5.0. This suggests that visual modalities are highly effective in correcting these types of errors.
@@ Appendix C Details about LLM-based Evaluation
Content Type Precision Recall F1
TERM 98.339 91.926 95.024
NUM 98.214 98.214 98.214
NE 70.000 100.000 82.353
GRAM 91.781 97.810 94.700
DISF 91.429 67.368 77.576
GEN 93.820 96.718 95.247
Micro Average 94.042 93.593 93.524
Severity Level Pearson Spearman
93.449 93.447
Table 6: The performance of LLM-based evaluator on Whisper-base transcripts of all videos, against gold labels annotated by humans. Pearson and Spearman correlation coefficient are calculated with three levels of severity converted into integers (0, 1, 2).

### v2

@@ Abstract
Recent advancements in multimodal large language models (MLLMs) have made significant progress in integrating information across various modalities, yet real-world applications in educational and scientific domains remain challenging. This paper introduces the Multimodal Scientific ASR (MS-ASR) task, which focuses on transcribing scientific conference videos by leveraging visual information from slides to enhance the accuracy of technical terminologies. Realized that traditional metrics like WER fall short in assessing performance accurately, prompting the proposal of severity-aware WER (SWER) that considers the content type and severity of ASR errors. We propose the Scientific Vision Augmented ASR (SciVASR) framework as a baseline method, enabling MLLMs to improve transcript quality through post-editing. Evaluations of state-of-the-art MLLMs, including GPT-4o, show a 45% improvement over speech-only baselines, highlighting the importance of multimodal information integration. [^1] Our code is available at https://github.com/yuriak/MS-ASR
@@
Minghan Wang1, Yuxia Wang2, Thuy-Trang Vu1, Ehsan Shareghi1, Gholamreza Haffari1 1Department of Data Science & AI, Monash University 2MBZUAI {minghan.wang,trang.vu1,ehsan.shareghi,gholamreza.haffari}@monash.edu yuxia.wang@mbzuai.ac.ae
@@ 1 Introduction
One such challenging task is multimodal automatic speech recognition (ASR) from knowledge-intensive presentation videos, which involves not only transcribing spoken content but also understanding and integrating visual information from images or videos (Oneata and Cucu, 2022; Chang et al., 2023). Presentation videos, often found in scientific conferences or educational contexts, contain dense, knowledge-rich content spanning various topics and concepts. Effective recognition and comprehension of such content are crucial for applications such as automatic transcript generation (Angrave et al., 2020), simultaneous translation (Wang et al., 2024b, a) in conference venue, lecture summarization (Benedetto et al., 2024) and accessibility tools for researchers and learners (McCarron, 2021). ASR on such videos presents several challenges, including keeping pace with the rapid evolution of technical domain knowledge and handling the variability in speaker accents, proficiency in English, and speech fluency (Cettolo et al., 2004). Mitigating these challenges requires ASR models to effectively extract and incorporate information from the speech and presentation slides in the video recordings (see Figure 1).
@@ 3.1 Speech Recognition
Figure 2: The architecture of SciVASR. Baseline ASR model transcribes a presentation recording into a transcript. Video segmenter splits video frames into a sequence of scenes with each presenting a slide. Slide analyzer applies multimodal LLMs to extract key information from scene images into visual contexts represented by text. Post-editor instruct a text-LLM to leverage visual contexts to refine the ASR transcript.
@@ 4 Experiments
Table 2: This table presents the full settings of our experiment showing the result in WER and Term-Recall (TR). Results are organized into 4 groups with various configurations (introduced in §4.2). Videos with $\star$ are from the test set. For clarity, we show the relative improvements compared to the Whisper-base with red and green distinguishing worse or better performance. Best results are highlighted with bold.
@@ 4.1 Models
ASR Models
We experiment with Whisper-base and Whisper-large-v3 (Radford et al., 2023) as speech recognition models. To challenge the effectiveness of our proposed method, we use the inferior Whisper-base as the baseline ASR whose outputs demands more complex refinement.
Vision-LLMs
We evaluate both open-sourced model CogAgent-VQA (Hong et al., 2023; Wang et al., 2023c) given the strengths in GUI and textual-related VQA tasks, and closed-sourced model GPT-4o (OpenAI et al., 2024).
Text-LLMs
Based on the observation that models with less than 13B parameters often fail to follow complex instructions and exhibit severe hallucinations, we apply open-sourced Llama3-70B-Instruct (AI@Meta, 2024) and closed-sourced GPT-4o  [^3] Model signature: gpt-4o-2024-05-13 in our experiments for effective error correction.
@@ 4.3 Results and Analysis
Why does post-editing increase WER?
Despite the significant improvements in term recall, WER doesn’t reveal consistent improvements. Open-sourced PE methods even increase the WER in 6 out of 10 videos. The discrepancy and the counter-intuition motivate us to investigate why the improvements aren’t reflected on WER. By analyzing the model outputs, we find that while post-editing accurately corrects a certain amount (roughly 40%) of explicit homophone errors, WER still judges them as errors due to the tiny differences with ground truth text in spelling, e.g., present vs. presents, and fine-tuning vs. finetuning. Additionally, acceptable minor rephrasing caused by the post-editing also increases WER, despite the improvements in fluency. These negligible errors are penalized similarly in WER compared with severe errors like recognizing BERT as birds, making WER fail to reflect the true quality of the transcript, significantly deviating from human judgments as shown in Figure 3. Post-edited transcripts are consistently voted to be better than the baseline transcript, and Vision-PE is more preferred than Text-PE results. To address this, we propose a severity-aware ASR evaluation protocol in §5.
Figure 3: The win, tie and lose ratio between different settings by human annotations, based on Whisper-base and SciVASR in open-source setup.
@@ 5.1 Breakdown of WER
WER is computed by counting the number of required operations (insertion, omission, substitution – $I,O,S$) to transform the hypothesis into the reference and then normalizing it by the reference length $N$. Each operation relates to a mismatch in the hypothesis. Formally, WER is calculated as $\frac{I+O+S}{N}$ where it assigns equal weights to all operations, without considering the varying importance and the semantic impact of the different types of mismatched content. This makes the phenomenon that a transcript with ten errors on general words like “the” or “a” has the same WER as a transcript with ten errors in terminologies. We mitigate it by re-weighting mismatches according to the content type and error severity.
@@
Content Type Category
@@ 5.2 ASR Error Taxonomy
Content Type Precision Recall F1
TERM 98.34 91.93 95.02
NUM 98.21 98.21 98.21
NE 70.00 100.00 82.35
GRAM 91.78 97.81 94.70
DISF 91.43 67.37 77.58
GEN 93.82 96.72 95.25
Micro Average 94.04 93.59 93.52
Severity Level Pearson Spearman
93.45 93.45
Table 4: The performance of LLM-based evaluator on Whisper-base transcripts of all videos, against gold labels annotated by humans. Pearson and Spearman correlation coefficients are calculated with three levels of severity converted into integers (0, 1, 2).
@@ 5.3 Severity-Aware WER
ASR Baselines
SciVASR with Open-sourced LLMs (Llama3-70B-Instruct + CogAgent-VQA)
SciVASR with Proprietary LLMs (GPT-4o)
@@
where $N$ is the reference length, $K$ is the number of mismatches which is same as the total number of operations required in WER, and $w_{i}$ is the weight corresponding to the severity level of each mismatch. In our experiment, we simply define the mapping of severity levels to weights as $w_{critical}$ = 1.0, $w_{minor}$ = 0.6, and $w_{ok}$ = 0.2. The rationale is to down-weight less severe mismatches and place greater emphasis on critical errors. We also tried the setting of (1.0, 0.5, 0.1) and found that (1.0, 0.6, 0.2) gives a better correlation with human evaluations on the dev set.
@@ 5.4 How reliable is LLM evaluation?
We prompt an LLM to label each mismatch with its content type category and severity level as shown in Figure 7 (see detailed prompt in Figure 6 in Appendix C). GPT-4o [^5] Model signature: gpt-4o-2024-05-13 is used in our experiments.
@@ 6 Re-Evaluating SciVASR
Compared with ASR baselines, text-only post-editing without visual information reduces both SWER and the content-wise severity scores to some extent. SWER decreases from 5.0 (Whisper-base) to 4.76 in open-sourced setting and 4.08 in close-sourced setting (Text-PE), indicating that parameterized knowledge in text-LLM is helpful in correcting potential errors.
Visual context is critical for correcting terminology, named entity and numbers
Incorporating visual information brings a significant reduction in the severity scores for TERM (-152), NE (-38) and NUM (-5) in the close-sourced group (GPT-4o Vision-PE vs Whisper-base). SWER reaches 2.75 from 5.0. This suggests that visual modalities are highly effective in correcting these types of errors.
@@ 7 Conclusion
Acknowledgments
The authors are grateful to the anonymous reviewers for their helpful comments and corrections. The first author was supported by an Australian Government Research Training Program (RTP) Scholarship. This work is supported by the ARC Future Fellowship FT190100039.